Convert the Polar form (2,-pie/6) to a complex number in the form a+ib.
Answer by scott8148(6628)

-pi/6 gives a 30º reference angle in quadrant IV

a = 2[cos(-30º)] = 2 * [(√3) / 2] = √3

b = 2[sin(-30º)] = 2 * (-1/2) = -1

√3 - i